QR is a diameter of a circle,centre O,and P is a point on its circumference.
Given that PQR= 25°, calculate
a, POR
b, OPR
c, the value, correct to 2 significant figures of PQ/QR.
1
Expert's answer
2020-07-19T15:54:05-0400

Consider the circle with center O and the diameter PQ


"central \ angle=intercepted\ arc"

The measure of each inscribed angle is exactly half the measure of its intercepted arc. 

In a circle, the measure of an inscribed angle is half the measure of the central angle with the same intercepted arc.

a. The angle "\angle PQR" is an inscribed angle


"m\angle PQR={1\over 2}m \overgroup{PR}"

The angle "\angle POR" is a central angle


"m\angle POR=m \overgroup{PR}=2m \angle PQR=2\cdot25\degree=50\degree"



b. We see that OP and OR are radii of the circle. Then "\Delta POR" is the equilateral triangle


"OP=OR, m\angle OPR=m\angle ORP"

The three interior angles in a triangle will always add up to 180°


"m\angle OPR+m\angle ORP+m\angle POR=180\degree"

"m\angle OPR+m\angle ORP=180\degree -m\angle POR"

"m\angle OPR=\dfrac{180\degree -m\angle POR}{2}"

"m\angle OPR=\dfrac{180\degree -50\degree}{2}=65\degree"

c. An angle inscribed in a semicircle is a right angle. Then


"m\angle QPR=90\degree"

Right "\Delta QPR"


"\cos\angle PQR=\dfrac{PQ}{QR}"

"\dfrac{PQ}{QR}=\cos25\degree\approx0.91"